OpenStudy (anonymous):

four less than twice the smaller number of two consecutive odd integers is eleven more than the larger odd integers i'm having trouble setting it up

OpenStudy (studygurl14):

"two consecutive odd integers" x and x + 2 "twice the small number" 2x "4 less" 2x - 4 "is" 2x - 4 = "larger interger" x + 2 "eleven more" x + 2 + 11 So... 2x - 4 = x + 2 + 11

OpenStudy (anonymous):

when i find the answer do i have to plug it in 2 find the other integer?

OpenStudy (studygurl14):

when you find x, solve "x + 2" to find the second one

OpenStudy (anonymous):

do i add 11 to the equation or not?

OpenStudy (studygurl14):

I don't understand what you mean... This is the equation to solve 2x - 4 = x + 2 + 11

OpenStudy (anonymous):

do i plug the answer to "x+2+11" or just "x+2"

OpenStudy (studygurl14):

just into "x+2"

OpenStudy (anonymous):

so the answer will be 19?

OpenStudy (anonymous):

thank you so much btw

OpenStudy (studygurl14):

The larger number would be 19, correct. You're welcome. Can I have a medal please? (click best answer)

OpenStudy (anonymous):

done and thank you again how do i find the smaller number?

OpenStudy (studygurl14):

The smaller number would be 19 - 2, which is 17

OpenStudy (anonymous):

okay thank so much

OpenStudy (studygurl14):

yw